# Supersaturation of Tree Posets

### v1

@@ 3. $(C_{2})$:
By $(C_{0})$ and $(C_{1})$ we have $\varphi^{j}(y)\in\mathcal{L}^{I(y)}(\mathcal{M}^{|P|-j})$. Because we constructed our marked chain families using Theorem 5.9, we have that $\varphi^{j}(y)$ is not $(I(y),\delta)$-lower bad with respect to $\mathcal{M}^{|P|-j-1}$. By Lemma 6.3, $\varphi^{j}(y)$ is not $(I(y),I(x),\gamma)$-bad with respect to $\mathcal{M}^{|P|-j-1}$. Suppose $|\mathcal{A}|<\gamma n^{\ell(I(y)-I(x))}$. We derive a contradiction by showing that $\varphi^{j}(y)$ is $(I(y),I(x),\gamma)$-bad with $\mathcal{W}_{1}=\mathcal{P}_{j}$ and $\mathcal{W}_{2}=\mathcal{A}$. Note that by definition of $\mathcal{P}_{0}=\mathcal{W}_{1}$ we have $\mathcal{W}_{1}\cap U(\varphi^{j}(y))=\emptyset$ and $|\mathcal{W}_{1}|\leq|P|$, and by hypothesis we are assuming $|\mathcal{W}_{2}|\leq\gamma n^{\ell|I(y)-I(x)|}$ It thus remains to check the last condition in the definition of $(I(y),I(x),\gamma)$-badness, i.e. that for each $(\chi,Q)\in\mathcal{M}(\varphi^{j}(y),I(y))$, either $Q\cap D^{*}(\varphi^{j}(y),\mathcal{W}_{1})\neq\emptyset$ or $Q$ contains a member from $\mathcal{W}_{2}$ in the $I(x)$th position.
Consider any $(\chi,Q)\in\mathcal{M}^{|P|-j-1}(\varphi^{j}(y),I(y))$. If $Q\cap D^{*}(\varphi^{j}(y),\mathcal{P}_{0})=\emptyset$, then the $I(x)$th member of $Q$ lies in $\mathcal{A}$ by definition of $\mathcal{A}$. Therefore, every $(\chi,Q)\in\mathcal{M}^{|P|-j-1}(\varphi^{j}(y),I(y))$ has $Q$ either intersecting $D^{*}(\varphi^{j}(y),\mathcal{P}_{j})$ or $\mathcal{A}$. Thus, $\varphi^{j}(y)$ is $(I(y),I(x),\gamma)$-bad, a contradiction. ∎
We now check that $\mathcal{A}$ is in fact a set of valid choices in the following sense. If we define $\varphi^{j+1}(v_{a})=\varphi^{j}(v_{a})$ for all $a\in[j]$ and $\varphi^{j+1}(x)=F$ for any $F\in\mathcal{A}$, then $\varphi^{j+1}$ maintains $(C_{1}),(C_{2})$. Since $\mathcal{M}^{|P|-j}\subseteq\mathcal{M}^{|P|-j-1}$, for all $a,b\in[j]$, $(C_{1})$ and $(C_{2})$ hold for all such pairs. Let $a\in[j]$, and suppose $v_{a}$ is not comparable with $x$. Then in particular $v_{a}\not\geq y$, so $\varphi^{j+1}(v_{a})=\varphi^{j}(v_{a})\in\mathcal{P}_{j}$. Since $F\in D(\varphi^{j}(y))$, yet $F\not\in D^{*}(\varphi^{j}(y),\mathcal{P}_{j})$, $F$ is not comparable with $\varphi^{j+1}(v_{a})$ in $\mathcal{B}_{n}$, thus $(C_{2})$ holds for all noncomparable pairs.
Note that by definition, $\varphi^{j+1}(x)$ is in the $I(x)$th position of some $(\chi,Q)\in\mathcal{M}^{|P|-j-1}$ which has $\varphi^{j+1}(y)$ in the $I(y)$th position. Thus, $(C_{1})$ holds for all edges in $H(P)[\{v_{1},\dots v_{j+1}\}]$, finishing the induction.
Let us now complete the calculation. In Case 1, we that the total number of ways is at least
$\Omega\left({n\choose n/2}\cdot\prod_{xy\in H(P)}n^{I(y)-I(x)}\right),$
giving the result. For Case 2, we instead have that the number of choices is at least
$\Omega\left(n^{-2}{n\choose n/2}\cdot\prod_{xy\in H(P)}n^{2(I(y)-I(x))}\right)\geq\Omega\left(n^{-2}{n\choose n/2}\cdot n^{|P|-1}\prod_{xy\in H(P)}n^{(I(y)-I(x))}\right),$
where in the first step we use the fact our family is $2$-gapped and the last step used that $I(y)\geq I(x)+1$ for each $xy\in H(P)$ since $I$ is admissible and that there are $|P|-1$ terms in the product. This also gives our desired count since we assumed $|P|\geq 3$, proving the result.
Lastly, we will need the following upper bound on $M(n,q,P)$, which we recall is the number of copies of $P$ in the $q$ middle levels of $\mathcal{B}_{n}$.
@@
If $P$ be a tree poset and $q$ is a positive integer, then
$M(n,q,P)\leq\sum_{I}\prod_{xy\in H(P)}n^{|I(x)-I(y)|}\cdot{n\choose n/2},$
where the sum ranges over all admissible $I:P\to[q]$ and where $H(P)$ denotes the Hasse diagram of $P$.
@@
Given a function $I:P\to[q]$, we say that a copy $P^{\prime}$ of $P$ in $\mathcal{M}_{n,q}$ is of type $I$ if for each $x\in P$, the member $F_{x}\in P^{\prime}$ corresponding to $x$ lies in the $I(x)$th level of $\mathcal{M}_{n,q}$ (here we emphasize that we mean $I(x)$th level from the top, e.g. the 1st level of $\mathcal{M}_{n,q}$ is the level whose member $F$ have the largest cardinality). Observe that every copy of $P$ must be of type $I$ for some admissible $I$, since if we had $I(x)\leq I(y)$ for $x<y$ then this would imply $F_{x}\geq F_{y}$ in $P^{\prime}$, a contradiction to $P^{\prime}$ being a copy of $P$ with $F_{x},F_{y}$ corresponding to $x,y$. As such, to prove the result it will suffice to show that for all admissible $I$, the number of copies of $P$ of type $I$ is at most
$\prod_{xy\in H(P)}n^{|I(x)-I(y)|}\cdot{n\choose n/2}.$
Let $(x_{1},\ldots,x_{|P|})$ be an ordering of the members of $P$ such that for each $x_{j}$ with $j>1$, there exists a unique $1\leq i<j$ such that $x_{i}x_{j}$ or $x_{j}x_{i}$ is an edge of $H(P)$. Note that such an ordering exists since $P$ is a tree poset. We can identify the copies of $P$ in $\mathcal{M}_{n,q}$ of type $I$ by tuples $(F_{1},\ldots,F_{|P|})$ of members of $\mathcal{M}_{n,q}$ such that the map $f(x_{i})=F_{i}$ defines a copy of $P$ of type $I$, so it will be enough to upper bound the number of such tuples.
Note that $F_{1}$ must be some members of the $I(x_{1})$th layer of $\mathcal{M}_{n,q}$, the number of which is at most (${n\choose n/2}$. Iteratively given that $F_{1},\ldots,F_{j-1}$ have been selected, we let $i$ be the unique $1\leq i<j$ such that $x_{i}x_{j}$ or $x_{j}x_{i}$ is an edge of $H(P)$. If $x_{i}<x_{j}$, then $F_{j}$ must be a set containing $F_{i}$ together with $I(x_{i})-I(x_{j})$ additional elements from $[n]$, and the number of such sets is at most $n^{I(x_{i})-I(x_{j})}$ (note that if $I(x_{i})\leq I(x_{j})$, then there exist no copies of $P$ of type $I$, so this upper bound continues to hold trivially). Similarly if $x_{j}<x_{i}$ then the number of sets is at most $n^{I(x_{j})-I(x_{i})}$. Multiplying the number of choices for each step gives the total result after observing that each edge of $H(P)$ is counted exactly once by some $n^{I(x_{i})-I(x_{j})}$ term. ∎

### v2

@@ 3. $(C_{3})$:
Let $v$ be any vertex of $P$ and $\gamma=\gamma(\varepsilon,q,\ell)$ the real number returned by Theorem 5.2 applied with $\varepsilon,q,\ell$.
Assume on the contrary that $\mathcal{F}$ contains less than
$\left(\frac{\gamma}{2}\right)^{|P|-1}\left(\prod_{xy\in H(P)}n^{\ell|\mathrm{r}(y)-\mathrm{r}(x)|}\right)\cdot N,$
induced copies of $P$.
Let $\mathcal{F}_{\mathrm{b}}$ be the set of elements in $\mathcal{F}$ which are embedded as $v$ in more than $\left(\frac{\gamma}{2}\right)^{|P|-1}\prod_{xy\in H(P)}n^{\ell|\mathrm{r}(y)-\mathrm{r}(x)|}$ induced copies of $P$.
Since by assumption, we know that
$|\mathcal{F}_{\mathrm{b}}|\cdot\left(\frac{\gamma}{2}\right)^{|P|-1}\prod_{xy\in H(P)}n^{\ell|\mathrm{r}(y)-\mathrm{r}(x)|}\leq\left(\frac{\gamma}{2}\right)^{|P|-1}\left(\prod_{xy\in H(P)}n^{\ell|\mathrm{r}(y)-\mathrm{r}(x)|}\right)\cdot N,$
$\implies|\mathcal{F}_{\mathrm{b}}|\leq N.$
Let $\mathcal{F}^{\prime}=\mathcal{F}-\mathcal{F}_{\mathrm{b}}$. Then, $\mu(\mathcal{F}^{\prime})\geq q-1+\varepsilon$ by assumption.
Now, apply Theorem 5.2 to $\mathcal{F}^{\prime}$ with $\Gamma=\emptyset,q,\ell,\varepsilon$. Then, there is $F\in\mathcal{F}$ playing the role of $v$ in at least
$\left(\frac{\gamma}{2}\right)^{|P|-1}\prod_{xy\in H(P)}n^{\ell|\mathrm{r}(y)-\mathrm{r}(x)|},$
induced copies of $P$. This contradicts the construction of $\mathcal{F}^{\prime}$, and so $\mathcal{F}$ contains at least
$\left(\frac{\gamma}{2}\right)^{|P|-1}\left(\prod_{xy\in H(P)}n^{\ell|\mathrm{r}(y)-\mathrm{r}(x)|}\right)\cdot N,$
induced copies of $P$, completing the proof of the theorem.
Lastly, we will need the following upper bound on $M^{*}(n,q,P)$ in terms of all poset homomorphisms $\mathrm{r}:P\rightarrow[q]$. Recall that $M^{*}(n,q,P)$ is the number of copies of $P$ in the $q$ middle levels of $\mathcal{B}_{n}$.
@@
If $P$ is a tree poset and $q$ is any positive integer, then
$M^{*}(n,q,P)\leq\sum_{\mathrm{r}}\prod_{xy\in H(P)}n^{|\mathrm{r}(x)-\mathrm{r}(y)|}\cdot{n\choose n/2},$
where the sum ranges over all poset homomorphisms $\mathrm{r}:P\to[q]$.
@@
For ease of notation, let $\mathcal{M}_{n,q}$ be the middle $q$ layers of $\mathcal{B}_{n}$. Given a poset homomorphism $\mathrm{r}:P\to[q]$, we say that a copy $P^{\prime}$ of $P$ in $\mathcal{M}_{n,q}$ is of type-$r$ if for each $x\in P$, the member $F_{x}\in P^{\prime}$ corresponding to $x$ lies in the $\mathrm{r}(x)$th level of $\mathcal{M}_{n,q}$. Observe that every copy of $P\in\mathcal{M}_{n,q}$ must be of type $\mathrm{r}$ for some $\mathrm{r}$. As such, to prove the result it will suffice to show that for all homomorphisms $\mathrm{r}$, the number of induced copies of $P$ or type $\mathrm{r}$ in $\mathcal{M}_{n,q}$ is at most
$\prod_{xy\in H(P)}n^{|\mathrm{r}(x)-\mathrm{r}(y)|}\cdot{n\choose n/2}.$
Let $x_{1},\ldots,x_{|P|}$ be an ordering of $P$ such that for all $j\geq 2$ $x_{j}$ has a unique neighbour, called the parent of $x_{j}$, among $x_{1},\dots,x_{j-1}$ if we view $H(P)$ as an undirected graph. Note that such an ordering exists since $P$ is a tree poset. We can identify the copies of $P$ in $\mathcal{M}_{n,q}$ of type-$\mathrm{r}$ by tuples $(F_{1},\ldots,F_{|P|})$ of members of $\mathcal{M}_{n,q}$ such that the map $f(x_{i})=F_{i}$ defines a copy of $P$ of type $\mathrm{r}$. So, it is enough to upper bound the number of such tuples.
Since $F_{1}$ must be a member of the $\mathrm{r}(x_{1})$th layer of $\mathcal{M}_{n,q}$, the number of choices for $F_{1}$ is at most (${n\choose n/2}$. Given that $F_{1},\ldots,F_{j-1}$ have been selected, let $x_{i}$ be the parent of $x_{j}$. If $x_{i}<x_{j}$, then $F_{j}$ must be a set containing $F_{i}$ together with $\mathrm{r}(x_{i})-\mathrm{r}(x_{j})$ additional elements from $[n]$, and the number of such sets is at most $n^{\mathrm{r}(x_{i})-\mathrm{r}(x_{j})}$. Similarly if $x_{j}<x_{i}$ then the number of choices for $F_{j}$ is at most $n^{\mathrm{r}(x_{j})-\mathrm{r}(x_{i})}$. Multiplying the number of choices for each step gives the total result, since each edge of $H(P)$ is counted exactly once by some $n^{|\mathrm{r}(x_{i})-\mathrm{r}(x_{j})|}$ term. ∎